lim(x,y)-(0,0) sinx/y exists.


1
Expert's answer
2021-03-01T06:42:33-0500

Solution:

"\lim _{(x,y)\rightarrow(0,0)} \sin (\frac xy)"

Let's calculate this in two different ways.

Along "x=0"

"\lim _{(x,y)\rightarrow(0,y)} \sin (\frac 0y)=0"

Now, along "y=0"

"\lim _{(x,y)\rightarrow(x,0)} \sin (\frac x0)=\infty"

Since, both the limits are unequal, limit of given function does not exist.